OpenStudy (anonymous):

solve for x: 4/x + 4/x^2-9 = 3/x-3

OpenStudy (anonymous):

Without any delimiters, ( ) or { } or [ ], your expression is:\[\frac{4}{x}+\frac{4}{x^2}-9=\frac{3}{x}-3 \]\[\left\{x=\frac{1}{12} \left(1-\sqrt{97}\right), x=\frac{1}{12} \left(1+\sqrt{97}\right)\right\}\]